Shift report 20010804 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Aug 04 2001 - 08:03:44 EDT

  • Next message: Apache: "Shift report 20010804 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Aug  4 08:03:44 2001, Shift: 00:00-08:00
    Supervisor: KH, with: A. Hanssen
    Log book#7 pages 17 - 18
    
    Shift Summary:
    ------------------------------------------------------------
    No beam for entire shift.
    
    Accelerator apparently recovered from fire.  Beam was 
    estimated at ~2:00am, but failed to materialize during 
    the whole night.
    
    In anticipation of beam at 2:00am, we
    Magnets were left on.
    
    TPC and DC anodes are off (except T1 anode is at 900V)
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Aug 04 2001 - 08:04:46 EDT